OpenStudy (anonymous):

2 sqrt108/ sqrt180y can you explain how to simplify this??

OpenStudy (anonymous):

\[\frac{2\sqrt{108}}{\sqrt{108y}}\] is that right?

OpenStudy (raden):

setting 108 = 36 * 3 180 = 36 * 5

OpenStudy (raden):

hmm.. why there is a y, that's not beautiful :)

OpenStudy (anonymous):

haha yea what do you do after that

OpenStudy (raden):

2 sqrt108/ sqrt180y = 2 sqrt(36*3) / sqrt(36*5*y) = 2 sqrt(36)sqrt(3) / sqrt(36)sqrt(5y) = 2 * 6 sqrt(3) / 6sqrt(5y) = 12sqrt(3)/6sqrt(5y) = (12/6) sqrt(3)/sqrt(5y) = 2 sqrt(3)/sqrt(5y) then u have to rationalize it

OpenStudy (anonymous):

ahh thanks

OpenStudy (raden):

welcome ;)